A hot brass plate is having its upper surface cooled by impinging jet of the air at temperature of 15 oC and convection coefficient of k=180, 190, 200, 210 220 and 230 W/m2 oK. The 20-cm wide and 10-cm length brass plate (ρ=8530 kg/m3, and Cp=380 J/kg K. k=220 w/mK had a uniform initial temperature of 650 oC, and lover surface of the plat is well insulated. Using uniform node for Δx = Δy = 2.5 cm and left side and right side of the plate cooled by natural convection of h=35 W/m2K and ambient temperature of 20 oC. Derive finite differential equation for all nodes at steady state.

We are suppose to write the equations for the 16 nodes that form in the grid, if someone could help me just showing the nodes, showing if I can apply similarity and writing a couple node equations so I can at least see how to start the problem, that would help. Thank you! I will like your work.

Air at 15°C h = 220 W/m2K L = 10-cm Insulation w = 20-cm
Answer #1

The detailed solution is given below.

Top 15°C hq=-220W/m%% 6 7 8 9 1 2 3 0 13 14 is 16 14 es D 0 22 21 23 24 25 24 27 1235W/mik To = 20°c 20 1 B D 19 h₂35w the To

Terminology let An=oy = a *درة 14) - ذر - Ag АкhЭ*1 An ya , = = 9 Р Ay- Диx1 . 1 г Д д Thus, An= Ay= on=oy=a

Mode #1 и 0 h, Съю, - т )+ h,, Сте, -т) Е 2. + кө, C-) + кву Стрь-т)=0 оо Node #2 2. 3 ыс h, 4, СТ», -ъ) + KA, Cта --) + КА»

lo 11 Nocle #lo 1 p h₂ An (Tooz-Tio) + KAY (Ti-To) + Karya (Tig- Tia) + KANCTU-Zoe o sy Equations are for mode # 19 8 28 are

are similar Equations for mode #23 6 32 Node #37 28 KAy CT2g-T37) + KAM CT3-737 AY than (Tø2-T37) =0 sn 37 38 2 Node #38 Insu

Node #41 4T41 zo 32 -040 bondaly Culing 2732 +2740 mirol for incubated & Symmetry boundary 40 3 2

Please upvote the solution if found satisfactory.